OpenStudy (anonymous):

Write an expression that represents the width of a rectangle with length x+5 and area x^3+3x^2+47x+60 A.x^2+17x+132+720/x+5 B.x^2+7x+12 C. x^3+7x^2+12x

OpenStudy (anonymous):

length*width=area so (x+5)*width=x^3+3x^2+47x+60 or width=(x^3+3x^2+47x+60)/(x+5)=? Sorry I'm not finding the answer; is there a typo or an answer D?